OpenStudy (anonymous):

Given sec theta= 4 lies in quadrant I. Find cos theta/2

OpenStudy (gorv):

cos and sec positve in first Quad

OpenStudy (gorv):

\[\theta= \sec ^{-1} 4\]

OpenStudy (gorv):

\[\cos \frac{ \theta }{ 2 }=\cos(\frac{ \sec ^{-1}4 }{ 2 })\]